1. DEF Company is comparing three different capital structures. Plan I would result in 800 shares of stock and $9,000 in debt. Plan II would result in 700 shares of stock and $13,500 in debt. Plan III is an all-equity plan and would result in 1,000 shares of stock. The firm’s EBIT will be $8,000 per year until infinity. The interest rate on the debt is 10%.                                                                                            (12 marks total)

a. Ignoring taxes, compute the EPS for each of the three plans. Which of the three plans has the highest EPS? Which has the lowest?                            (4 marks)

b. Compute the break-even EBIT that will cause the EPS on Plan I to be equal to the all-equity EPS.                                                                                       (2 marks)

c.   Compute the break-even EBIT that will cause the EPS on Plan II to be equal to the all-equity EPS.                                                                         (2 marks)

d. Compare your results from parts (b) and (c) above. Is one higher than the other? Why?                                                                                                          (1 mark)

e. Ignoring taxes, what is the break-even EBIT that will cause the EPS on Plan I to be equal to the EPS on Plan II? What conclusions do you reach when you compare the outcomes of parts (b), (c), and (e) above?                  (3 marks)
Answer #1

As per rules I am answering the first 4 subparts of the question

EPS = Net Income/Shares outstanding

A: Plan I EPS = (8000-10%*9000)/800

8.875

Plan II EPS= (8000-10%*13500)/700

=9.5

Plan III EPS = 8000/1000 = 8

Plan II is highest and Plan III is lowest.

B: Let Breakeven EBIT be E

(E-10%*9000)/800 = E/1000

1000E-900000= 800E

200E= 900,000

E = 4500

C: Let Breakeven EBIT be E

(E-10%*13500)/700 = E/1000

1000E-1350,000= 700E

300E = 1350,000

E = 4500

D: Both are equal due to different interest amounts and shares outstanding.